# Solving the Correlation Cluster LP in Sublinear Time

### v2

@@ Abstract
[11] introduced the cluster LP for Correlation Clustering, which they argued captures the problem much more succinctly than previous linear programming formulations. However, the cluster LP has exponential size, with a variable for every possible set of vertices in the input graph. Nevertheless, [11] showed how to find a feasible solution for the cluster LP in time $O(n^{\text{poly}(1/\varepsilon)})$ with objective value at most $(1+\varepsilon)$ times the value of an optimal solution for the respective Correlation Clustering instance. Furthermore, they showed how to round a solution to the cluster LP, yielding a $(1.437+\varepsilon)$-approximation algorithm for the Correlation Clustering problem.
The main technical result of this paper is a new approach to find a feasible solution for the cluster LP with objective value at most $(1+\varepsilon)$ of the optimum in time $\widetilde{O}(2^{\text{poly}(1/\varepsilon)}n)$, where $n$ is the number of vertices in the graph. We also show how to implement the rounding within the same time bounds, thus achieving a fast $(1.437+\varepsilon)$-approximation algorithm for the Correlation Clustering problem. This bridges the gap between state-of-the-art methods for approximating Correlation Clustering and the recent focus on fast algorithms.
@@ 1 Introduction
Recent work has shown how to surpass this barrier. Cohen-Addad, Lee and Newman employed the Sherali-Adams hierarchy to obtain a $(1.994+\varepsilon)$-approximation [17], which was later improved to $(1.73+\varepsilon)$ by Cohen-Addad, Lee, Li and Newman, using a new so-called preclustering technique to preprocess the instance [15]. Most recently, Cao, Cohen-Addad, Lee, Li, Newman and Vogl introduced the cluster LP framework which generalizes all previously known formulations, and showed how to round it to obtain a $(1.437+\varepsilon)$-approximation in polynomial time [11]. However, the cluster LP (see Section 1.1 for details) has exponential size: It contains a variable for each subset of the vertices indicating whether this subset is a cluster or not. The recent work of [11] leverages the preclustering method to compute a solution to the cluster LP in time $O(n^{\text{poly}(1/\varepsilon)})$ whose value is within a $(1+\varepsilon)$-factor of the cost an optimal clustering, hence leading to a polynomial-time approximation algorithm.
@@
In general, the approximation guarantees obtained in these various restricted settings have remained significantly higher than the best known polynomial time approximation of $(1.437+\varepsilon)$, which employed computationally expensive solutions to Sherali-Adams relaxations of Correlation Clustering to solve the cluster LP.
Thus, it is natural to consider what we can achieve when we are allowed limited computation time. For instance, How well can Correlation Clustering be approximated in (sub)linear time? Here, by linear time, we mean time linear in the number of $+$edges, and by sublinear time, we mean time linear in the number of vertices.
@@
Before we explain our results in more detail, we introduce some basic notation. The input to the correlation clustering problem is a complete graph where each edge is labeled either as a $+$edge or a --edge. For a graph $G$, we often denote its vertex set by $V(G)$ and its edge set by $E(G)$. We let $G=(V,E)$ represent the subgraph induced by the $+$edges (i.e., $E^{+}=E(G)$), while the set of --edges is given by $E^{-}=\binom{V(G)}{2}\setminus E(G)$. Let $n=|V|$ and $m=|E|$ represent the number of vertices and $+$edges in $G$, respectively. Given a graph $G$ and a subset $V^{\prime}\subseteq V(G)$, $G[V^{\prime}]$ denotes the subgraph of $G$ induced by $V^{\prime}$. For technical reasons, we assume $E$ contains all $n$ self-loops $\{uu:u\in V\}$.
@@ 1.1 Our Results
Let $\varepsilon,\delta>0$ be small enough constants and let $\mathrm{OPT}$ be the cost of the optimum solution to the given Correlation Clustering instance. Then there is a small $\Delta={\mathrm{poly}}(\varepsilon)$ such that the following statement holds. One can output a solution $(z_{S})_{S\subseteq V}$ to the cluster LP with $\mathrm{obj}(x)\leq(1+\varepsilon)\mathrm{OPT}$ in expectation, described using a list of non-zero coordinates such that each coordinate of $z$ is either 00 or at least $\Delta$. In the various models, the respective procedure has the following attributes.
@@
Let $\varepsilon>0$ be a small enough constant and $\Delta={\mathrm{poly}}(1/\varepsilon)$. Given a solution to the cluster LP $(z_{S})_{S\subseteq V}$ where each coordinate of $z$ is either 00 or at least $\Delta$, one can output a clustering with expected cost at most $(1.437+\varepsilon)\mathrm{obj}(x)$ in time $\tilde{O}(n/\Delta^{2})$.
@@
There exists a $(1.437+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $\tilde{O}(2^{{\mathrm{poly}}(1/\varepsilon)}n)$.
@@ 1.2 Technical Overview
Given a set $S$, let $E^{+}(S,V\setminus S)=\{uv\in E^{+}\mid u\in S,v\not\in S\textmd{ or }u\not\in S,v\in S\}$ denote the set of $+$edges with exactly one endpoint in $S$, and let $E^{-}(S)=\{uv\in E^{-}\mid u,v\in S\}$ denote the set of --edges with both endpoints in $S$. We can rewrite the objective value of cluster LP using $z$ values instead of $x$ values as
@@
[…]
@@
$\max$ $\sum_{u\in V}q_{u}\leavevmode\nobreak$ (dual cluster LP)
@@ 2 Preclustering
We first introduce some more necessary notation. Define $d^{+}(v)$ as the degree of vertex $v$ with respect to $+$edges in $G$. For any subset $C\subseteq V$, let $d^{+}(v,C)$ and $d^{-}(v,C)$ denote the number of $+$edges and --edges from $v$ to $C$, respectively, connecting $v$ to vertices in $C$. When the context is clear, we omit the $+$ symbol, so $d(v)$ and $d(v,C)$ mean $d^{+}(v)$ and $d^{+}(v,C)$, respectively.
@@
One can think of this model as storing an adjacency list and an adjacency matrix of $+$edges for $G$. The matrix allows us to check whether $uv\in E^{+}$ in $O(1)$ time.
@@ 2.2 Preclustering
For a vertex $u$, let $d_{\mathrm{adm}}(u)$ denote the number of vertices $v\in V$ such that $(u,v)$ is an admissible pair. Note that admissible, atomic, and non-admissible edges can each be either $+$edges or --edges. Let $N_{\mathrm{adm}}(v)$ to be the set of vertices $u$ such that $(u,v)\in E_{\mathrm{adm}}$. We use $K(v)$ to represent the set $K\in{\mathcal{K}}$ that contains $v$; to make the notation more general, when $v$ is a singleton atom, $K(v)=\{v\}$.
@@
  • (MPC model) The algorithm takes $O(1)$ rounds succeeds with probability at least $1-1/n$ and requires $O(n^{\delta})$ memory per machine. Moreover, the algorithm uses a total memory of $O(m\log n)$, where $m$ is the number of the $+$ edges.
@@ 3.1 Converting cluster LP to covering cluster LP
[…]
@@
Then, we will prove the lower bound on $d_{\mathrm{cross}}(K)$. Consider a vertex $v\in K$ and an admissible edge $(v,u)$. If $uv$ is a $+$edge, then $d_{\mathrm{cross}}(K)$ already covers it, so we only need to consider --edges.
Let $A(v)$ be the set of all vertices that are connected to $v$ by an admissible --edge. Let $u$ be one such neighbor in $A(v)$. By Definition 5, $v$ and $u$ are degree similar and must share at least $\varepsilon\cdot\min\{d(v),d(u)\}\geq\varepsilon^{2}d(v)$ $+$neighbors, which are degree similar to both $u$ and $v$. Recall that a pair $w\tilde{w}$ is degree similar if $\varepsilon d(w)\leq d(\tilde{w})\leq d(w)/\varepsilon$. We will distinguish between two cases.
Let $A_{1}(v)\subseteq A(v)$ be the set of vertices $u\in A(v)$ such that at least half of the degree-similar $+$neighbors of $u$ and $v$ are not in $K$. Note that $v\in K$, has at most $d(v)-d(v,K)$ degree similar $+$neighbors outside $K$. Each $u\in A_{1}(v)$ has to be adjacent to at least $\frac{\varepsilon^{2}}{2}d(v)$ of these vertices outside of $K$ which have degree at most $d(v)/\varepsilon$. Therefore, we have at most
@@
Let $A_{2}(v)\subseteq A(v)$ be the set of vertices $u\in A(v)$ such that that at least half of the degree-similar $+$neighbors of $u$ and $v$ are in $K$. Each $u\in A_{2}(v)$ must connect to at least $\frac{\varepsilon^{2}}{2}d(v)$ vertices in $K$, so we have at most
@@
Combining the two cases, we find that the --admissible neighbors of $K$ are at most $\sum_{v\in K}|A(v)|=O(\varepsilon^{-3}d_{\mathrm{cross}}(K))$. ∎
@@
which is true since the --edges in $U$ also belong to $W$ and $+$edges contributing to $\mathrm{cost}(U)$ but not $\mathrm{cost}(W)$ are at most those in $E^{+}(U,W\setminus{U})$, whose contribution to $\mathrm{cost}(U)$ is $|E^{+}(U,W\setminus{U})|/2<d_{\mathrm{cross}}(W\setminus{U})$. ∎
@@
$\mathrm{cost}(\widetilde{z})$ $={\mathrm{cover}}(\widetilde{z})-d_{\mathrm{cross}}(V)\leq{\mathrm{cover}}(z)-d_{\mathrm{cross}}(V)$
@@ 3.2 The MWU Algorithm
  • has objective value ${\mathrm{cover}}(z^{(t)})\leq(1+5\gamma)\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$,
@@
Let $\varepsilon>0$ be a small enough constant and $\gamma=O(\varepsilon^{13})$. After $T_{\mathrm{MW}}=\frac{\log(1/\gamma)}{\gamma^{4}}$ rounds, Algorithm 2 returns a solution $z$ to the covering cluster LP with objective ${\mathrm{cover}}(z)\leq(1+O(\gamma))\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$.
@@
$\sum_{v\in U}d_{\mathrm{cross}}(v)\leq 2\gamma\leavevmode\nobreak\ d_{\mathrm{cross}}(V).$
@@
Next, we will bound the objective value of $z^{\star}$. By the second property of Lemma 15, we have ${\mathrm{cover}}(z^{(t)})\leq(1+5\gamma)\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$ for each $t\leq T_{\mathrm{MW}}$. Since ${\mathrm{cover}}$ is a linear function, this upper bound also holds for the average $\frac{1}{T_{\mathrm{MW}}}\sum_{t=1}^{T_{\mathrm{MW}}}z^{(t)}$. At the end of Algorithm 2, we cover all uncovered atoms $U=\{v\in V\mid\sum_{S\ni v}\hat{z}_{S}\leq 1-2\gamma\}$. When we set a coordinate $\hat{z}_{K}$ to $1$, the ${\mathrm{cover}}$ objective value increases by at most $3/2\cdot d_{\mathrm{cross}}(K)$. We have $\mathrm{cost}(K)=1/2\cdot d_{\mathrm{cross}}(K)$. The additional $d_{\mathrm{cross}}(K)$ is the term added to $\mathrm{cost}(K)$ for the covering cluster LP. Recall that each $z^{(t)}$ does not split atoms. Hence, if $v\in U$ then $K(v)\subseteq U$. By Claim 14, the cost increases by at most,
@@ 4 Finding a Partial Clustering with Small Ratio in Polynomial Time
    has objective ${\mathrm{cover}}(z)\leq(1+5\gamma)\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})$,
@@
In order to construct the point $z$, we will find a partial clustering =$\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ that achieves a small ratio $\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}$. Each vertex $v$ is contained in at most one set $S\in\mathcal{F}$. Moreover, $\mathcal{F}$ is a partial clustering, meaning that $\mathcal{F}$ might not cover all the vertices in $V$. However, we require that $\mathcal{F}$ covers at least a constant fraction of the probability mass of the vertices (i.e., $p(\mathcal{F})\geq\gamma$). We show that Algorithm 3 will find such a partial clustering.
@@
8:     Add $C$ to $\mathcal{F}$, set $\hat{p}_{v}$ to 00 for all $v\in C$ and remove vertices in $C$ from $\hat{V}$.
@@
Given vertex weights $p_{v}>0$ for all $v\in V$, Algorithm 3 finds a family =$\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $S,T\in\mathcal{F}$, $S\cap T=\emptyset$,
@@
First, observe that Property 3. holds by the condition of the while loop. Property 4. holds since the all small ratio clusters $\{C_{v}\}_{v\in V}$ do not split atoms by Lemma 18. Next, we observe that Property 1. holds: the family $\mathcal{F}$ consists of disjoint sets. Observe that after we added a cluster $C\subseteq V$ to $\mathcal{F}$ we set the weight $\hat{p}_{v}$ to 00 for all $v\in C$. Note that if the weight $\hat{p}_{v}$ is set to 00 it remains 00 throughout the execution of the algorithm. Moreover, when we add a cluster $C$ to $\mathcal{F}$, it contains only vertices such that $\hat{p}_{v}>0$. Thus, when we add $C$ to $\mathcal{F}$ we have $\hat{p}_{v}>0$ for all $v\in C$ and $\hat{p}_{v}=0$ for all $v\in S,S\in\mathcal{F}$.
@@
For the second inequality, we use that $\gamma\leavevmode\nobreak\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\leavevmode\nobreak\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). The last inequality holds by the definition of $C^{\star}$. Let $v\in C^{\star}$ such that if $K$ is a non-singleton atom in $C^{\star}$ then $K=K(v)$. By (11),
@@
Finding the small ratio cluster $C_{v}$ for a single vertex $v$ takes time $\widetilde{O}(d^{2}(v))$. So overall, we need time $\widetilde{O}(n^{3})$ to find all $\{C_{v}\}_{v\in V}$. There are at most $n$ iterations of the while loop on Line 6 since we set $\hat{p}_{v}=0$ for at least one vertex $v$. This can happen at most $n$ times. Indeed, once a vertex has weight $\hat{p}_{v}=0$, the weight will remain 00. During one iteration of the while loop on Line 6, we might have to update all $\{C_{v}\}_{v\in V}$. Again, this takes time at most $\widetilde{O}(n^{3})$. ∎

### v5

@@ Abstract
[11] introduced the cluster LP for Correlation Clustering, which they argued captures the problem much more succinctly than previous linear programming formulations. However, the cluster LP has exponential size, with a variable for every possible set of vertices in the input graph. Nevertheless, [11] showed how to find a feasible solution for the cluster LP in time $O(n^{\text{poly}(1/\varepsilon)})$ with objective value at most $(1+\varepsilon)$ times the value of an optimal solution for the respective Correlation Clustering instance. Furthermore, they showed how to round a solution to the cluster LP, yielding a $(1.485+\varepsilon)$-approximation algorithm for the Correlation Clustering problem.111[11] claimed that there is a rounding for the cluster LP yielding a 1.437-approximation algorithm for Correlation Clustering. However, the proof has a bug, and the best bound we can currently prove for the cluster LP is 1.485. Proof details will appear shortly in an updated arxiv version of [11].
The main technical result of this paper is a new approach to find a feasible solution for the cluster LP with objective value at most $(1+\varepsilon)$ of the optimum in time $\widetilde{O}(2^{\text{poly}(1/\varepsilon)}n)$, where $n$ is the number of vertices in the graph. We also show how to implement the rounding within the same time bounds, thus achieving a fast $(1.485+\varepsilon)$-approximation algorithm for the Correlation Clustering problem. This bridges the gap between state-of-the-art methods for approximating Correlation Clustering and the recent focus on fast algorithms.
@@ 1 Introduction
Recent work has shown how to surpass this barrier. Cohen-Addad, Lee and Newman employed the Sherali-Adams hierarchy to obtain a $(1.994+\varepsilon)$-approximation [17], which was later improved to $(1.73+\varepsilon)$ by Cohen-Addad, Lee, Li and Newman, using a new so-called preclustering technique to preprocess the instance [15]. Most recently, Cao, Cohen-Addad, Lee, Li, Newman and Vogl introduced the cluster LP framework which generalizes all previously known formulations, and showed how to round it to obtain a $(1.485+\varepsilon)$-approximation in polynomial time [11].  [^2] See the footnote in the abstract for a remark regarding the 1.485 approximation ratio for cluster LP. However, the cluster LP (see Section 1.1 for details) has exponential size: It contains a variable for each subset of the vertices indicating whether this subset is a cluster or not. The recent work of [11] leverages the preclustering method to compute a solution to the cluster LP in time $O(n^{\text{poly}(1/\varepsilon)})$ whose value is within a $(1+\varepsilon)$-factor of the cost an optimal clustering, hence leading to a polynomial-time approximation algorithm.
@@
In general, the approximation guarantees obtained in these various restricted settings have remained significantly higher than the best known polynomial time approximation of $(1.485+\varepsilon)$, which employed computationally expensive solutions to Sherali-Adams relaxations of Correlation Clustering to solve the cluster LP.
Thus, it is natural to consider what we can achieve when we are allowed limited computation time. For instance, How well can Correlation Clustering be approximated in (sub)linear time? Here, by linear time, we mean time linear in the number of ++edges, and by sublinear time, we mean time linear in the number of vertices.
@@
Before we explain our results in more detail, we introduce some basic notation. The input to the correlation clustering problem is a complete graph where each edge is labeled either as a ++edge or a -edge. For a graph $G$, we often denote its vertex set by $V(G)$ and its edge set by $E(G)$. We let $G=(V,E)$ represent the subgraph induced by the ++edges (i.e., $E^{+}=E(G)$), while the set of -edges is given by $E^{-}=\binom{V(G)}{2}\setminus E(G)$. Let $n=|V|$ and $m=|E|$ represent the number of vertices and ++edges in $G$, respectively. Given a graph $G$ and a subset $V^{\prime}\subseteq V(G)$, $G[V^{\prime}]$ denotes the subgraph of $G$ induced by $V^{\prime}$. For technical reasons, we assume $E$ contains all $n$ self-loops $\{uu:u\in V\}$.
@@ 1.1 Our Results
Let $\varepsilon,\delta>0$ be small enough constants and let $\mathrm{OPT}$ be the cost of the optimum solution to the given Correlation Clustering instance. Then there is a small $\Delta={\mathrm{poly}}(\varepsilon)$ such that the following statement holds. One can output a solution $(z_{S})_{S\subseteq V}$ to the cluster LP with $\mathrm{obj}(x)\leq(1+\varepsilon)\mathrm{OPT}$ in expectation, described using a list of non-zero coordinates such that each coordinate of $z$ is either 0 or at least $\Delta$. In the various models, the respective procedure has the following attributes.
@@
Let $\varepsilon>0$ be a small enough constant and $\Delta={\mathrm{poly}}(1/\varepsilon)$. Given a solution to the cluster LP $(z_{S})_{S\subseteq V}$ where each coordinate of $z$ is either 0 or at least $\Delta$, one can output a clustering with expected cost at most $(1.485+\varepsilon)\mathrm{obj}(x)$ in time $\tilde{O}(n/\Delta^{2})$.
@@
There exists a $(1.485+\varepsilon)$-approximation algorithm for Correlation Clustering that runs in time $\tilde{O}(2^{{\mathrm{poly}}(1/\varepsilon)}n)$.
@@ 1.2 Technical Overview
Given a set $S$, let $E^{+}(S,V\setminus S)=\{uv\in E^{+}\mid u\in S,v\not\in S\textmd{ or }u\not\in S,v\in S\}$ denote the set of ++edges with exactly one endpoint in $S$, and let $E^{-}(S)=\{uv\in E^{-}\mid u,v\in S\}$ denote the set of -edges with both endpoints in $S$. We can rewrite the objective value of cluster LP using $z$ values instead of $x$ values as
@@
$\min\qquad\mathrm{cost}(z)\qquad\text{s.t.}$ (cluster LP)
@@
$\max$ $\sum_{u\in V}q_{u}\penalty 10000\$(dual cluster LP)
@@ 2 Preclustering
We first introduce some more necessary notation. Define $d^{+}(v)$ as the degree of vertex $v$ with respect to ++edges in $G$. For any subset $C\subseteq V$, let $d^{+}(v,C)$ and $d^{-}(v,C)$ denote the number of ++edges and -edges from $v$ to $C$, respectively, connecting $v$ to vertices in $C$. When the context is clear, we omit the ++ symbol, so $d(v)$ and $d(v,C)$ mean $d^{+}(v)$ and $d^{+}(v,C)$, respectively.
@@
One can think of this model as storing an adjacency list and an adjacency matrix of ++edges for $G$. The matrix allows us to check whether $uv\in E^{+}$ in $O(1)$ time.
@@ 2.2 Preclustering
For a vertex $u$, let $d_{\mathrm{adm}}(u)$ denote the number of vertices $v\in V$ such that $(u,v)$ is an admissible pair. Note that admissible, atomic, and non-admissible edges can each be either ++edges or -edges. Let $N_{\mathrm{adm}}(v)$ to be the set of vertices $u$ such that $(u,v)\in E_{\mathrm{adm}}$. We use $K(v)$ to represent the set $K\in{\mathcal{K}}$ that contains $v$; to make the notation more general, when $v$ is a singleton atom, $K(v)=\{v\}$.
@@
  • (MPC model) The algorithm takes $O(1)$ rounds succeeds with probability at least 11/n1-1/n and requires $O(n^{\delta})$ memory per machine. Moreover, the algorithm uses a total memory of $O(m\log n)$, where $m$ is the number of the ++ edges.
@@ 3.1 Converting cluster LP to covering cluster LP
$\min\qquad{\mathrm{cover}}(z)\qquad\text{s.t.}$ (covering cluster LP)
@@
Then, we will prove the lower bound on $d_{\mathrm{cross}}(K)$. Consider a vertex $v\in K$ and an admissible edge $(v,u)$. If $uv$ is a ++edge, then $d_{\mathrm{cross}}(K)$ already covers it, so we only need to consider -edges.
Let $A(v)$ be the set of all vertices that are connected to $v$ by an admissible -edge. Let $u$ be one such neighbor in $A(v)$. By Definition 5, $v$ and $u$ are degree similar and must share at least $\varepsilon\cdot\min\{d(v),d(u)\}\geq\varepsilon^{2}d(v)$ ++neighbors, which are degree similar to both $u$ and $v$. Recall that a pair $w\tilde{w}$ is degree similar if $\varepsilon d(w)\leq d(\tilde{w})\leq d(w)/\varepsilon$. We will distinguish between two cases.
Let $A_{1}(v)\subseteq A(v)$ be the set of vertices $u\in A(v)$ such that at least half of the degree-similar ++neighbors of $u$ and $v$ are not in $K$. Note that $v\in K$, has at most d(v)d(v,K)d(v)-d(v,K) degree similar ++neighbors outside $K$. Each $u\in A_{1}(v)$ has to be adjacent to at least $\frac{\varepsilon^{2}}{2}d(v)$ of these vertices outside of $K$ which have degree at most $d(v)/\varepsilon$. Therefore, we have at most
@@
Let $A_{2}(v)\subseteq A(v)$ be the set of vertices $u\in A(v)$ such that that at least half of the degree-similar ++neighbors of $u$ and $v$ are in $K$. Each $u\in A_{2}(v)$ must connect to at least $\frac{\varepsilon^{2}}{2}d(v)$ vertices in $K$, so we have at most
@@
Combining the two cases, we find that the -admissible neighbors of $K$ are at most $\sum_{v\in K}|A(v)|=O(\varepsilon^{-3}d_{\mathrm{cross}}(K))$. ∎
@@
which is true since the -edges in $U$ also belong to $W$ and ++edges contributing to $\mathrm{cost}(U)$ but not $\mathrm{cost}(W)$ are at most those in $E^{+}(U,W\setminus{U})$, whose contribution to $\mathrm{cost}(U)$ is $|E^{+}(U,W\setminus{U})|/2<d_{\mathrm{cross}}(W\setminus{U})$. ∎
@@
[…]
@@ 3.2 The MWU Algorithm
  • has objective value ${\mathrm{cover}}(z^{(t)})\leq(1+5\gamma)\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$,
@@
Let $\varepsilon>0$ be a small enough constant and $\gamma=O(\varepsilon^{13})$. After $T_{\mathrm{MW}}=\frac{\log(1/\gamma)}{\gamma^{4}}$ rounds, Algorithm 2 returns a solution $z$ to the covering cluster LP with objective ${\mathrm{cover}}(z)\leq(1+O(\gamma))\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$.
@@
$\sum_{v\in U}d_{\mathrm{cross}}(v)\leq 2\gamma\penalty 10000\ d_{\mathrm{cross}}(V).$
@@
Next, we will bound the objective value of $z^{\star}$. By the second property of Lemma 15, we have ${\mathrm{cover}}(z^{(t)})\leq(1+5\gamma)\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$ for each $t\leq T_{\mathrm{MW}}$. Since $\mathrm{cover}$} is a linear function, this upper bound also holds for the average $\frac{1}{T_{\mathrm{MW}}}\sum_{t=1}^{T_{\mathrm{MW}}}z^{(t)}$. At the end of Algorithm 2, we cover all uncovered atoms $U=\{v\in V\mid\sum_{S\ni v}\hat{z}_{S}\leq 1-2\gamma\}$. When we set a coordinate $\hat{z}_{K}$ to $1$, the $\mathrm{cover}$} objective value increases by at most $3/2\cdot d_{\mathrm{cross}}(K)$. We have $\mathrm{cost}(K)=1/2\cdot d_{\mathrm{cross}}(K)$. The additional $d_{\mathrm{cross}}(K)$ is the term added to $\mathrm{cost}(K)$ for the covering cluster LP. Recall that each $z^{(t)}$ does not split atoms. Hence, if $v\in U$ then $K(v)\subseteq U$. By Claim 14, the cost increases by at most,
@@ 4 Finding a Partial Clustering with Small Ratio in Polynomial Time
    has objective ${\mathrm{cover}}(z)\leq(1+5\gamma)\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})$,
@@
In order to construct the point $z$, we will find a partial clustering $\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ that achieves a small ratio $\frac{{\mathrm{cover}}(\mathcal{F})}{p(\mathcal{F})}$. Each vertex $v$ is contained in at most one set $S\in\mathcal{F}$. Moreover, $\mathcal{F}$ is a partial clustering, meaning that $\mathcal{F}$ might not cover all the vertices in $V$. However, we require that $\mathcal{F}$ covers at least a constant fraction of the probability mass of the vertices (i.e., $p(\mathcal{F})\geq\gamma$). We show that Algorithm 3 will find such a partial clustering.
@@
8:  Add $C$ to $\mathcal{F}$, set $\hat{p}_{v}$ to 0 for all $v\in C$ and remove vertices in $C$ from $\hat{V}$.
@@
Given vertex weights $p_{v}>0$ for all $v\in V$, Algorithm 3 finds a family $\mathcal{F}=\{S_{1},S_{2},...,S_{l}\mid S_{i}\subseteq V\}$ such that,
@@
    for any distinct $\in\mathcal{F}, ST=S\cap T$$\emptyset$,
@@
First, observe that Property 3. holds by the condition of the while loop. Property 4. holds since the all small ratio clusters $\{C_{v}\}_{v\in V}$ do not split atoms by Lemma 18. Next, we observe that Property 1. holds: the family $\mathcal{F}$ consists of disjoint sets. Observe that after we added a cluster $C\subseteq V$ to $\mathcal{F}$ we set the weight $\hat{p}_{v}$ to 0 for all $v\in C$. Note that if the weight $\hat{p}_{v}$ is set to 0 it remains 0 throughout the execution of the algorithm. Moreover, when we add a cluster $C$ to $\mathcal{F}$, it contains only vertices such that $\hat{p}_{v}>0$. Thus, when we add $C$ to $\mathcal{F}$ we have $\hat{p}_{v}>0$ for all $v\in C$ and $\hat{p}_{v}=0$ for all $v\in S,S\in\mathcal{F}$.
@@
For the second inequality, we use that $\gamma\penalty 10000\ {\mathrm{cover}}(\mathrm{OPT})\geq\gamma\penalty 10000\ \mathrm{cost}(\mathrm{OPT})\geq\gamma^{2}|E_{\text{adm}}|$ because of the preclustering (Theorem 7). The last inequality holds by the definition of $C^{\star}$. Let $v\in C^{\star}$ such that if $K$ is a non-singleton atom in $C^{\star}$ then $K=K(v)$. By (11),
@@
Finding the small ratio cluster $C_{v}$ for a single vertex $v$ takes time $\widetilde{O}(d^{2}(v))$. So overall, we need time $\widetilde{O}(n^{3})$ to find all $\{C_{v}\}_{v\in V}$. There are at most $n$ iterations of the while loop on Line 6 since we set $\hat{p}_{v}=0$ for at least one vertex $v$. This can happen at most $n$ times. Indeed, once a vertex has weight $\hat{p}_{v}=0$, the weight will remain 0. During one iteration of the while loop on Line 6, we might have to update all $\{C_{v}\}_{v\in V}$. Again, this takes time at most $\widetilde{O}(n^{3})$. ∎